OpenStudy (anonymous):

Maggie has a job working in an office for $9 an hour and another job driving a tractor for $18 an hour. Maggie works at both jobs and earns $630 in one week. (a) Write an equation that gives the number of hours, x, she worked for $9 per hour and the number of hours, y, she worked for $18 per hour. (b) Find the x- and y-intercepts for this equation. x-intercept (x, y) y-intercept (x, y)

OpenStudy (anonymous):

@kelliegirl33 do you have any idea ?

OpenStudy (anonymous):

9x + 18y = 630 x intercept : 9x + 18(0) = 630 x = 630/9 x = 70 <-- x intercept y intercept : 9(0) + 18y = 630 18y = 630 y = 630/18 y = 35 <-- y intercept 9x + 18y = 630 9(70) + 18(35) = 630 not sure about the hours for each......

OpenStudy (anonymous):

x intercept = (70,0) y intercept = (0,35)

OpenStudy (anonymous):

thank you ive been struggling with these problems for weeks

OpenStudy (anonymous):

then it asks me this (d) From the graph, find how many hours she worked at $18 if she worked 38 hours at $9 per hour. hr (e) From the graph, find how many hours she worked at $9 if she worked 22 hours at $18 per hour. hr